What is the area of a circle with a diameter of 20 yards ?
Answer by lenny460(1073)

What is the area of a circle with a diameter of 20 yards ?
Area of a circle = Pi(r^2)
Pi = 3.14
Diameter = 20 yards.
Therefore:
Radius = 10 yards.
3.14 * 10^2
= 3.14 * 100 = 314


Answer:
314 yards squared










Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com